# An Analysis of Decision Problems for Relational Pattern Languages under Various Constraints

### v1

@@ 0.B.2 Definition of the Universal Turing Machine $U$
for every configuration $(q_{i},t_{L},t_{R})$. Recall that $i>0$ as $q_{i}\in\{q_{1},...,q_{15}\}$. A computation $\mathcal{C}=(C_{1},...,C_{n})$ on $U$ is a finite sequence of configurations of $U$. It is valid if $C_{1}=I$ (II being some initial configuration), $C_{n}$ is a halting configuration, and $C_{i+1}$ is a valid successor configuration of $C_{i}$, for $i\in[n-1]$, as defined by $\delta$. In [BREMER201215], the notion is adopted that any possible configuration where both tape sides have a finite value under $e$ is a valid successor configuration of a halting configuration. The encoding of computations of $U$ is given analogously to the definition in the case of nondeterministic 2-counter automata without input, i.e., for some computation $\mathcal{C}=(C_{1},...,c_{n})$, we have
@@ 0 $(0,R,q_{1})$ $(1,L,q_{11})$ $(0,R,q_{12})$ $(0,R,q_{13})$ $(0,L,q_{2})$ $(0,L,q_{3})$ $(0,R,q_{14})$
Table 4: Transition table of $U$, i.e., definition of $\delta$, as it is given in [BREMER201215] or [Neary2009].
0.B.3 Undecidability of the Inclusion Problem for Several Basic Relations
Proof of Corollary˜1:
Proof
Consider the construction of Freydenberger and Reidenbach [Freydenberger2010] and the construction of Bremer and Freydenberger [BREMER201215], respectively. The main differences between E-pattern languages $L_{E}(\alpha)$ (resp. NE pattern languages $L_{NE}(\alpha)$) of a pattern $\alpha\in Pat_{\Sigma}$ in [Freydenberger2010, BREMER201215]and relational pattern languages $L_{E}(\alpha,r_{\alpha})$ (resp. $L_{NE}(\alpha,r_{\alpha})$) that we consider are
  1. 1.
    In relational patterns each variable may occur only once whereas in classical patterns each variable may occur any number of times.
  2. 2.
    In relational pattern languages the substitution must be $r_{\alpha}$-valid whereas in pattern languages there is no relation.
Thus, if we use the pattern given in relational pattern languages to simulate identical variables, there is no significant difference between relational pattern languages and classical pattern languages. With this, we can immediately refer to the proofs in [Freydenberger2010, BREMER201215]. First, consider $r_{=}$. We can use the proofs in [Freydenberger2010, BREMER201215]. For each variable $x$, we replace each multiple occurrence of $x$ by new variables $x_{i},i\in\mathbb{N}$ with $(r_{=},x,x_{i})\in r_{\alpha}$. Since $(x,x_{i})\in r_{\alpha}$, we get $h(x)=h(x_{i})$ for each $r_{\alpha}$-valid substitution $h$. As $x_{i}$ is a new variable, we do not change anything else. Thus, this slightly adapted proof works in exactly the same way as the original proof and we get that the inclusion problem for relational pattern languages with the relation $r_{=}$ is undecidable.
Now, we show that we can do the same with the relations $r_{ssq}$ and $r_{*}$. For $r_{ssq}$, we get $(r_{ssq},x,y),(r_{ssq},y,x)\in r_{\alpha}$ implies $h(x)=h(y)$ for any two variables $x,y$ and any $r_{\alpha}$-valid substitution $h$ by the definition of $r_{ssq}$. For each variable $x$, we replace each multiple occurrence of $x$ by new variables $x_{i},i\in\mathbb{N}$ with $(r_{ssq},x,x_{i}),(r_{ssq},x_{i},x)\in r_{\alpha}$. With this, the proof works analogously to the one of [Freydenberger2010, BREMER201215]. For $r_{*}$, we get the same observation: $(r_{*},x,y),(r_{*},y,x)\in r_{\alpha}$ implies $h(x)=h(y)$ for any two variables $x,y$ and any $r_{\alpha}$-valid substitution $h$. Thus, the proof works here as well.
So, in total, we get that the inclusion problem for relational pattern languages with the relation $\tilde{r}$ for $\tilde{r}\in\{r_{=},r_{ssq},r_{*}\}$ is undecidable.
0.B.4 Undecidability of the Inclusion Problem for Non-Erasing Relational Pattern Languages under Abelian Equivalence
Proof
As this proof works almost identical to the proof of Bremer and Freydenberer in [BREMER201215], we only give an extended sketch and focus on the formal intricacies that occur as we are now considering abelian equivalence instead of the equivalence relation.
Let $\Sigma=\{0,\#\}$ and $R:=\{r_{ab}\}$. To show undecidability of this problem, we reduce the problem of deciding whether $\mathtt{UnivValc}_{NE}(I)=\emptyset$ for some initial configuration $I$ of $U$ to the problem of deciding whether $L_{NE}(p_{1},\alpha_{R})\subseteq L_{NE}(p_{2},\beta_{R})$ for two relational patterns $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$. To be specific, we construct two such patterns $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$ such that $L_{NE}(\alpha,r_{\alpha})\subseteq L_{NE}(\beta,r_{\beta})$ if and only if $\mathtt{ValC}_{U}(I)=\emptyset$.
@@ 0.B.3 Undecidability of the Inclusion Problem for Several Basic Relations
where $v:=0\#^{4}0$, $\alpha_{1}$ is some relational pattern not containing $^{3}$ as a factor, $\alpha_{2}$ is a relational pattern not containing #\#, and $t$ is some other terminal string. As we shall see later, we have $r_{\alpha}=\emptyset$. This will become clear once we specifically define $\alpha_{1}$ and $\alpha_{2}$. We define $t$ using the non-erasing substitution $\psi:(\operatorname{var}(\beta)\cup\Sigma)^{*}\rightarrow\Sigma^{*}$ with $\psi(x)=0$ for all $x\in\operatorname{var}(\beta)$. Using that, we say $t:=\psi(r_{1}\hat{\beta}_{1}\cdots r_{\mu}\hat{\beta}_{\mu}r_{\mu+1})$. We see that both patterns are basically constructed in the same way as in [BREMER201215], just using the formalism of relational patterns, i.e. all variables which are equal in the construction in [BREMER201215], are in this construction distinct variables that are, however, related to each other through the sets $r_{\alpha}$ and $r_{\beta}$.
@@
The following lemma follows directly from the proof in [BREMER201215] which can be converted directly to the formalism using substitutions over relational patterns.
Lemma 6
@@
[BREMER201215] All $\psi(\hat{\beta}_{i})$ with $1\leq i\leq\mu$ and $t$ begin and end with the terminal letter 0 and do not contain the terminal word $^{4}$ as a factor.
@@
We recall, that $H_{r_{\alpha}}$ denotes the set of all $r_{\alpha}$-valid substitutions. Let $H_{r_{\alpha}^{+}}$ denote the subset of all non-erasing substitutions of $H_{r_{\alpha}}$. Each pair $(\upgamma_{i},\updelta_{i})$ is interpreted as a predicate $\pi_{i}:H_{r_{\alpha}^{+}}\rightarrow\{0,1\}$ such that any $h\in H_{r_{\alpha}^{+}}$ satisfies $\pi_{i}$ if there exists some $r_{\beta}$-valid non-erasing substitution $\tau:(\operatorname{var}(\upgamma_{i}\updelta_{i})\cap\Sigma)^{*}\rightarrow\Sigma^{*}$ with $\tau(\upgamma_{i})=h(0\alpha_{1}0)$ and $\tau(\updelta_{i})=h(0\alpha_{2}0)$. We need to say that $\tau$ is $r_{\beta}$-valid for formal reasons, but as all variables in $\updelta_{i}\upgamma_{i}$ are only related to variables in $\updelta_{i}\upgamma_{i}$ itself, we can consider such an isolated substitution as it does not interfere with the substitution of any other variable of $\beta$. Using those predicates, we will see that $L_{NE}(\alpha,r_{\alpha})\setminus L_{NE}(\beta,r_{\beta})$ contains those words $h(\alpha)$ for $h\in H_{r_{\alpha}^{+}}$ for which we have that $h$ does not satisfy any predicate $\pi_{i}$. Essentially, all predicates $\pi_{1}$ to $\pi_{\mu}$ will give an exhaustive list of criteria for membership of $\overline{\mathtt{ValC}_{U}(I)}$. We proceed with more technical preperations that we need for the easy and dynamic construction of predicates. Once this is settled, as the constructed predicates from [BREMER201215] also work in this case, we are done.
@@
We say that a substitution $h\in H_{r_{\alpha}^{+}}$ is of bad form if $h(0\alpha_{1}0)$ contains a factor $^{3}$ or if $h(0\alpha_{2}0)$ contains the letter #\#. We use the predicates $\pi_{1}$ and $\pi_{2}$ to handle all such substitutions of bad form and construct them exactly the same way as they are constructed in [BREMER201215] by saying
@@
for new variables $y_{1,1},y_{1,2},y_{2},\hat{y}_{1},\hat{y}_{2,1},\hat{y}_{2,2}\in X$ that are not related to any other variable in $\beta$. As they are not related to any other variable, no intricacies regarding abelian equivalence in comparison to normal equality occur and we directly get the following lemma from [BREMER201215] as their proof needs no adaptation.
Lemma 7
[BREMER201215] A substitution $h\in H_{r_{\alpha}^{+}}$ is of bad-form if and only if $h$ satisfies $\pi_{1}$ or $\pi_{2}$.
@@
Proof
Due to the different formalism, we need to rewrite certain parts of the proof from [BREMER201215]. However, the general idea is the same. We just have to show that it also works for abelian equivalence.
@@
Notice that none of the variables in one factor $\operatorname{var}(\upgamma_{i}\updelta_{i})$ appear outside of $\upgamma_{i}$ and $\updelta_{i}$. By that, we can always define $\tau^{\prime}$ that way. Also note, that the only significant difference to [BREMER201215] lies in the definition of $\tau^{\prime}(x_{j,k})$, $\tau^{\prime}(a_{k})$, and $\tau^{\prime}(b_{k})$ as no variables can occur multiple timesin relational patterns; the rest remains the same.
As in [BREMER201215] we obtain $\tau^{\prime}(\upgamma_{i})=\tau(\gamma_{i})=h(0\alpha_{1}0)$, $\tau^{\prime}(\updelta_{i})=\tau(\updelta_{i})=h(0\alpha_{1}0)$, and
@@
As these equations must hold and especially as $h(0\alpha_{1}0)=\tau(\upgamma_{j})$ and $h(0\alpha_{2}0)=\tau(\updelta_{j})$, we know that $h$ satisfies the predicate $\pi_{j}$. ∎
@@
For technical reasons, we define a predicate that sets a lower bound to the length of $h(\alpha_{2})$ if unsatisfied for a substitution $h\in H_{r_{\alpha}^{+}}$ of good form. As in [BREMER201215], the predicate $\pi_{3}$ is defined by
@@
for pairwise non-related variables $y_{3,1},...,y_{3,4}$ and $\hat{y}_{3,1},...,\hat{3,3}$. As in [BREMER201215], if some $h\in H_{r_{\alpha}^{+}}$ satisfies $\pi_{3}$, then $h(\alpha_{2})$ is a concatenation of three non-empty factors of $h(\alpha_{1})$. In other words, if some $h$ does not satisfy any of $\pi_{1}$$\pi_{2}$, and $\pi_{3}$, then $h(\alpha_{2})\in 0^{+}$ has to be longer than the $3$ longest non-overlapping and non-touching factors of 0’s in $h(\alpha_{1})$. This restriction allows for a simpler way to construct all other predicates $\pi_{4}$ to $\pi_{\mu}$.
@@
As we will see, the next part regarding simple predicates is also very similar to the one defined and used in [BREMER201215] but had to be adapted formally to also work in the context of relational patterns. Let $X_{s}:=\{\hat{x}_{1,1},\hat{x}_{2,1},\hat{x}_{3,1},\hat{x}_{1,2},\hat{x}_{2,2},\hat{x}_{3,2},...\}\subset X$ and always assume that $(r_{ab},\hat{x}_{i,j},\hat{x}_{i,j^{\prime}})\in r_{p}$ for $i\in\{1,2,3\}$ and $j,j^{\prime}\in\mathbb{N}$ in any set of relational constraints $r_{p}$ of any relational pattern $(p,r_{p})\in\mathtt{RelPat}_{\Sigma,R}$ if any variable of $X_{s}$ is used in $p$. Let $G_{r_{\alpha}^{+}}\subset H_{r_{\alpha}^{+}}$ denote the subset of $r_{\alpha}$-valid substitutions that are of good form and let $S$ be the set of all non-erasing substitutions $s:(\Sigma\cup X_{s})^{*}\rightarrow\Sigma^{*}$ for which $s(\hat{x}_{i,j})\in\{0\}^{+}$, for all $i\in\{1,2,3\}$ and $j\in\mathbb{N}$, and for which we have $s(\hat{x}_{i,j})=s(\hat{x}_{i^{\prime},j^{\prime}})$ if $i=i^{\prime}$ for $i,i^{\prime}\in\{1,2,3\}$ and $j,j^{\prime}\in\mathbb{N}$. For any relational pattern $(p_{s},r_{p_{s}})$ with $p_{s}\in(\Sigma\cup X_{s})^{*}\cap\mathtt{RelPat}_{\Sigma,R}$ we define $S(p_{s},r_{p_{s}}):=\{s(p_{s})|s\in S\}$. The next definition is exactly the same as in [BREMER201215], just adapted to the notation of relational patterns.
Definition 1
@@
[BREMER201215] A predicate $\pi:G_{r_{\alpha}^{+}}\rightarrow\{0,1\}$ is called a NE-simple predicate for $0\alpha_{1}0$ if there exists a relational pattern $(p_{s},r_{p_{s}})$ with $p_{s}\in(\Sigma\cup X_{s})^{*}\cap\mathtt{RelPat}_{\Sigma,R}$ and $(r_{ab},\hat{x}_{i,j},\hat{x}_{i,j^{\prime}})\in r_{p_{s}}$ for $i\in\{1,2,3\}$ and $j,j^{\prime}\in\mathbb{N}$ and languages $L_{1}\in\{0\Sigma^{*},\{0\}\}$ and $L_{2}\in\{\Sigma^{*}0,\{0\}\}$ such that any non-erasing substitution $h\in H_{r_{\alpha}^{+}}$ satisfies $\pi$ if and only if $h(0\alpha_{1}0)\in L_{1}S(p_{s},r_{p_{s}})L_{2}$. If $L_{1}=0\Sigma^{*}$ and $L_{2}=\Sigma^{*}0$, then we call $\pi$ an infix-predicate. If $L_{1}=\{0\}$ and $L_{2}=\Sigma^{*}0$, then we call $\pi$ a prefix-predicate in if its the other way around, we call $\pi$ a suffix predicate.
@@
We understand elements of $X_{s}$ as numerical paramters that describe (concatenative) powers of 0 with non-erasing substitutions $s:(\Sigma\cup X_{s})^{*}\rightarrow\Sigma^{*}$ acting as assignments. The next lemma is the final step we need to take in order to finish the adaptation of the original proof to the new case. It shows that for each simple predicate $\pi_{s}$, we can construct a predicate $\pi$ defined over some $\upgamma$ and $\updelta$ such that each substitution of good form that would have satisfied $\pi_{s}$ now satisfies either $\pi_{3}$ or $\pi_{s}$ which suffices the demands of the construction. As we can see, the argument is again very similar to the one in [BREMER201215] and had to be adapted mainly to the new formalism.
Lemma 9
@@
The other cases, where $L_{1}$ or $L_{2}$ is chosen to be the language containing only the word 0, are handled analogously by omitting $y_{1}$ or $y_{2}$ and keeping the rest of the construction. ∎

### v2

@@ Appendix 0.B Additional Content for the Inclusion Problem
for every configuration $(q_{i},t_{L},t_{R})$. Recall that $i>0$ as $q_{i}\in\{q_{1},...,q_{15}\}$. A computation $\mathcal{C}=(C_{1},...,C_{n})$ on $U$ is a finite sequence of configurations of $U$. It is valid if $C_{1}=I$ (II being some initial configuration), $C_{n}$ is a halting configuration, and $C_{i+1}$ is a valid successor configuration of $C_{i}$, for $i\in[n-1]$, as defined by $\delta$. In [5], the notion is adopted that any possible configuration where both tape sides have a finite value under $e$ is a valid successor configuration of a halting configuration. The encoding of computations of $U$ is given analogously to the definition in the case of nondeterministic 2-counter automata without input, i.e., for some computation $\mathcal{C}=(C_{1},...,c_{n})$, we have
@@ 0 $(0,R,q_{1})$ $(1,L,q_{11})$ $(0,R,q_{12})$ $(0,R,q_{13})$ $(0,L,q_{2})$ $(0,L,q_{3})$ $(0,R,q_{14})$
Table 2: Transition table of $U$, i.e., definition of $\delta$, as it is given in [5] or [31].
0.B.1.2 (The Full Proof Idea):
Proof.
As this proof works almost identical to the proof of Bremer and Freydenberer in [5], we only give an extended sketch and focus on the formal intricacies that occur as we are now considering abelian equivalence instead of the equivalence relation.
Let $\Sigma=\{0,\#\}$ and w.l.o.g. $R:=\{r_{ab}\}$. The arguments work analogously for $r_{rev}$ as both, $r_{rev}$ and $r_{ab}$, force equality on unary words and, in particular, single letters.
To show undecidability of this problem, we reduce the problem of deciding whether $\mathtt{UnivValc}_{NE}(I)=\emptyset$ for some initial configuration $I$ of $U$ to the problem of deciding whether $L_{NE}(p_{1},\alpha_{R})\subseteq L_{NE}(p_{2},\beta_{R})$ for two relational patterns $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$. To be specific, we construct two such patterns $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$ such that $L_{NE}(\alpha,r_{\alpha})\subseteq L_{NE}(\beta,r_{\beta})$ if and only if $\mathtt{ValC}_{U}(I)=\emptyset$.
@@
where $v:=0\#^{4}0$, $\alpha_{1}$ is some relational pattern not containing $^{3}$ as a factor, $\alpha_{2}$ is a relational pattern not containing #\#, and $t$ is some other terminal string. As we shall see later, we have $r_{\alpha}=\emptyset$. This will become clear once we specifically define $\alpha_{1}$ and $\alpha_{2}$. We define $t$ using the non-erasing substitution $\psi:(\operatorname{var}(\beta)\cup\Sigma)^{*}\rightarrow\Sigma^{*}$ with $\psi(x)=0$ for all $x\in\operatorname{var}(\beta)$. Using that, we say $t:=\psi(r_{1}\hat{\beta}_{1}\cdots r_{\mu}\hat{\beta}_{\mu}r_{\mu+1})$. We see that both patterns are basically constructed in the same way as in [5], just using the formalism of relational patterns, i.e. all variables which are equal in the construction in [5], are in this construction distinct variables that are, however, related to each other through the sets $r_{\alpha}$ and $r_{\beta}$.
@@
The following lemma follows directly from the proof in [5] which can be converted directly to the formalism using substitutions over relational patterns.
Lemma 6.
@@
[5] All $\psi(\hat{\beta}_{i})$ with $1\leq i\leq\mu$ and $t$ begin and end with the terminal letter 0 and do not contain the terminal word $^{4}$ as a factor.
@@
We recall, that $H_{r_{\alpha}}$ denotes the set of all $r_{\alpha}$-valid substitutions. Let $H_{r_{\alpha}^{+}}$ denote the subset of all non-erasing substitutions of $H_{r_{\alpha}}$. Each pair $(\upgamma_{i},\updelta_{i})$ is interpreted as a predicate $\pi_{i}:H_{r_{\alpha}^{+}}\rightarrow\{0,1\}$ such that any $h\in H_{r_{\alpha}^{+}}$ satisfies $\pi_{i}$ if there exists some $r_{\beta}$-valid non-erasing substitution $\tau:(\operatorname{var}(\upgamma_{i}\updelta_{i})\cap\Sigma)^{*}\rightarrow\Sigma^{*}$ with $\tau(\upgamma_{i})=h(0\alpha_{1}0)$ and $\tau(\updelta_{i})=h(0\alpha_{2}0)$. We need to say that $\tau$ is $r_{\beta}$-valid for formal reasons, but as all variables in $\updelta_{i}\upgamma_{i}$ are only related to variables in $\updelta_{i}\upgamma_{i}$ itself, we can consider such an isolated substitution as it does not interfere with the substitution of any other variable of $\beta$. Using those predicates, we will see that $L_{NE}(\alpha,r_{\alpha})\setminus L_{NE}(\beta,r_{\beta})$ contains those words $h(\alpha)$ for $h\in H_{r_{\alpha}^{+}}$ for which we have that $h$ does not satisfy any predicate $\pi_{i}$. Essentially, all predicates $\pi_{1}$ to $\pi_{\mu}$ will give an exhaustive list of criteria for membership of $\overline{\mathtt{ValC}_{U}(I)}$. We proceed with more technical preperations that we need for the easy and dynamic construction of predicates. Once this is settled, as the constructed predicates from [5] also work in this case, we are done.
@@
We say that a substitution $h\in H_{r_{\alpha}^{+}}$ is of bad form if $h(0\alpha_{1}0)$ contains a factor $^{3}$ or if $h(0\alpha_{2}0)$ contains the letter #\#. We use the predicates $\pi_{1}$ and $\pi_{2}$ to handle all such substitutions of bad form and construct them exactly the same way as they are constructed in [5] by saying
@@
for new variables $y_{1,1},y_{1,2},y_{2},\hat{y}_{1},\hat{y}_{2,1},\hat{y}_{2,2}\in X$ that are not related to any other variable in $\beta$. As they are not related to any other variable, no intricacies regarding abelian equivalence in comparison to normal equality occur and we directly get the following lemma from [5] as their proof needs no adaptation.
Lemma 7.
[5] A substitution $h\in H_{r_{\alpha}^{+}}$ is of bad-form if and only if $h$ satisfies $\pi_{1}$ or $\pi_{2}$.
@@
Proof.
Due to the different formalism, we need to rewrite certain parts of the proof from [5]. However, the general idea is the same. We just have to show that it also works for abelian equivalence.
@@
Notice that none of the variables in one factor $\operatorname{var}(\upgamma_{i}\updelta_{i})$ appear outside of $\upgamma_{i}$ and $\updelta_{i}$. By that, we can always define $\tau^{\prime}$ that way. Also note, that the only significant difference to [5] lies in the definition of $\tau^{\prime}(x_{j,k})$, $\tau^{\prime}(a_{k})$, and $\tau^{\prime}(b_{k})$ as no variables can occur multiple timesin relational patterns; the rest remains the same.
As in [5] we obtain $\tau^{\prime}(\upgamma_{i})=\tau(\gamma_{i})=h(0\alpha_{1}0)$, $\tau^{\prime}(\updelta_{i})=\tau(\updelta_{i})=h(0\alpha_{1}0)$, and
@@
As these equations must hold and especially as $h(0\alpha_{1}0)=\tau(\upgamma_{j})$ and $h(0\alpha_{2}0)=\tau(\updelta_{j})$, we know that $h$ satisfies the predicate $\pi_{j}$. ∎∎
@@
For technical reasons, we define a predicate that sets a lower bound to the length of $h(\alpha_{2})$ if unsatisfied for a substitution $h\in H_{r_{\alpha}^{+}}$ of good form. As in [5], the predicate $\pi_{3}$ is defined by
@@
for pairwise non-related variables $y_{3,1},...,y_{3,4}$ and $\hat{y}_{3,1},...,\hat{3,3}$. As in [5], if some $h\in H_{r_{\alpha}^{+}}$ satisfies $\pi_{3}$, then $h(\alpha_{2})$ is a concatenation of three non-empty factors of $h(\alpha_{1})$. In other words, if some $h$ does not satisfy any of $\pi_{1}$$\pi_{2}$, and $\pi_{3}$, then $h(\alpha_{2})\in 0^{+}$ has to be longer than the $3$ longest non-overlapping and non-touching factors of 0’s in $h(\alpha_{1})$. This restriction allows for a simpler way to construct all other predicates $\pi_{4}$ to $\pi_{\mu}$.
@@
As we will see, the next part regarding simple predicates is also very similar to the one defined and used in [5] but had to be adapted formally to also work in the context of relational patterns. Let $X_{s}:=\{\hat{x}_{1,1},\hat{x}_{2,1},\hat{x}_{3,1},\hat{x}_{1,2},\hat{x}_{2,2},\hat{x}_{3,2},...\}\subset X$ and always assume that $(r_{ab},\hat{x}_{i,j},\hat{x}_{i,j^{\prime}})\in r_{p}$ for $i\in\{1,2,3\}$ and $j,j^{\prime}\in\mathbb{N}$ in any set of relational constraints $r_{p}$ of any relational pattern $(p,r_{p})\in\mathtt{RelPat}_{\Sigma,R}$ if any variable of $X_{s}$ is used in $p$. Let $G_{r_{\alpha}^{+}}\subset H_{r_{\alpha}^{+}}$ denote the subset of $r_{\alpha}$-valid substitutions that are of good form and let $S$ be the set of all non-erasing substitutions $s:(\Sigma\cup X_{s})^{*}\rightarrow\Sigma^{*}$ for which $s(\hat{x}_{i,j})\in\{0\}^{+}$, for all $i\in\{1,2,3\}$ and $j\in\mathbb{N}$, and for which we have $s(\hat{x}_{i,j})=s(\hat{x}_{i^{\prime},j^{\prime}})$ if $i=i^{\prime}$ for $i,i^{\prime}\in\{1,2,3\}$ and $j,j^{\prime}\in\mathbb{N}$. For any relational pattern $(p_{s},r_{p_{s}})$ with $p_{s}\in(\Sigma\cup X_{s})^{*}\cap\mathtt{RelPat}_{\Sigma,R}$ we define $S(p_{s},r_{p_{s}}):=\{s(p_{s})|s\in S\}$. The next definition is exactly the same as in [5], just adapted to the notation of relational patterns.
Definition 1.
@@
[5] A predicate $\pi:G_{r_{\alpha}^{+}}\rightarrow\{0,1\}$ is called a NE-simple predicate for $0\alpha_{1}0$ if there exists a relational pattern $(p_{s},r_{p_{s}})$ with $p_{s}\in(\Sigma\cup X_{s})^{*}\cap\mathtt{RelPat}_{\Sigma,R}$ and $(r_{ab},\hat{x}_{i,j},\hat{x}_{i,j^{\prime}})\in r_{p_{s}}$ for $i\in\{1,2,3\}$ and $j,j^{\prime}\in\mathbb{N}$ and languages $L_{1}\in\{0\Sigma^{*},\{0\}\}$ and $L_{2}\in\{\Sigma^{*}0,\{0\}\}$ such that any non-erasing substitution $h\in H_{r_{\alpha}^{+}}$ satisfies $\pi$ if and only if $h(0\alpha_{1}0)\in L_{1}S(p_{s},r_{p_{s}})L_{2}$. If $L_{1}=0\Sigma^{*}$ and $L_{2}=\Sigma^{*}0$, then we call $\pi$ an infix-predicate. If $L_{1}=\{0\}$ and $L_{2}=\Sigma^{*}0$, then we call $\pi$ a prefix-predicate in if its the other way around, we call $\pi$ a suffix predicate.
@@
We understand elements of $X_{s}$ as numerical paramters that describe (concatenative) powers of 0 with non-erasing substitutions $s:(\Sigma\cup X_{s})^{*}\rightarrow\Sigma^{*}$ acting as assignments. The next lemma is the final step we need to take in order to finish the adaptation of the original proof to the new case. It shows that for each simple predicate $\pi_{s}$, we can construct a predicate $\pi$ defined over some $\upgamma$ and $\updelta$ such that each substitution of good form that would have satisfied $\pi_{s}$ now satisfies either $\pi_{3}$ or $\pi_{s}$ which suffices the demands of the construction. As we can see, the argument is again very similar to the one in [5] and had to be adapted mainly to the new formalism.
Lemma 9.
@@
The other cases, where $L_{1}$ or $L_{2}$ is chosen to be the language containing only the word 0, are handled analogously by omitting $y_{1}$ or $y_{2}$ and keeping the rest of the construction. ∎∎